OpenStudy (anonymous):

Prove sqrt((1+sinx)/(1-sinx)) = ((1+sinx)/(absolute value of cosx))

OpenStudy (anonymous):

hint: sin^2+cos^2=1

OpenStudy (akashdeepdeb):

\[\sqrt{\frac{(1+sinx)}{(1-sinx)}} = \sqrt{\frac{(1+sinx)*(1+ \sin x)}{(1-sinx)*(1+\sin x)}} = \sqrt{\frac{(1+sinx)^2}{(\cos^2 x)}} = {\frac{(1+sinx)}{|(cos x)|}}\] Got it? :)

OpenStudy (akashdeepdeb):

In the third part, I have used the first most basic identity in trigonometry sin^ x + cos^2 x = 1

OpenStudy (anonymous):

thanks

OpenStudy (akashdeepdeb):

:)